in how many years will a 73 year old grandfather be 5 times as old as his 5 year old grandson
Found 3 solutions by josgarithmetic, Solver92311, greenestamps:
Answer by josgarithmetic(39618)
Answer by Solver92311(821)





Solve for

John

My calculator said it, I believe it, that settles it

From
I > Ø

Answer by greenestamps(13200)



The difference between their ages is 73-5=68 years.

When the grandfather is 5 times as old as his grandson -- i.e., 4 times OLDER THAN -- the grandson's age will be 68/4=17 and the grandfather's age will be 17*5=85.

Since their current ages are 5 and 73, that will be 12 years from now.